OpenStudy (dejmccray):

I forgot how to do this and I'm being dumb If f(x) = 0.5(1 – x), what is the value of f(–1)? A.0.5 B.1 C.2.5 D.0

OpenStudy (ksaimouli):

Plug in x as -1

OpenStudy (texaschic101):

yep..its that simple f(x) = 0.5(1 - x) f(-1) = 0.5(1 -(-1) can you take it from here ?